OpenStudy (anonymous):

What is the electric force between a +3 µC point charge and a +6 µC point charge if they are separated by a distance of 0.75 m? (µC = 1.0 × 10^–6 C) ***How do I show the work for this? Thank you!!

OpenStudy (michele_laino):

here we have to apply the Coulomb law, namely the force of interaction, which is a repulsive force, is given by the subsequent formula: \[\Large F = K\frac{{{Q_1}{Q_2}}}{{{d^2}}}\]

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

i'm not sure what do i sub in? :/

OpenStudy (michele_laino):

where K=9*10^9 Q_1, Q_2 are the two electric charges, and d is their separation. So substituting, we have: \[\Large F = K\frac{{{Q_1}{Q_2}}}{{{d^2}}} = 9 \times {10^9}\frac{{3 \times {{10}^{ - 6}} \times 6 \times {{10}^{ - 6}}}}{{{{\left( {0.75} \right)}^2}}}\]

OpenStudy (anonymous):

ok! how would i insert that into my calc? would i get 0.288? i do not know if i inserted it correctly? :/

OpenStudy (michele_laino):

that's right! The correct result is a force whose modulus is: F=0.288 Newton

OpenStudy (anonymous):

oh yay!! :) is that the answer? or is there more? or is 0.288 newton the answer? that is the electrical force?

OpenStudy (michele_laino):

yes! We can draw that force as below: 

OpenStudy (anonymous):

ooh yay!! so we are done?

OpenStudy (michele_laino):

yes! we have finished!

OpenStudy (anonymous):

yay!! thank you1!

OpenStudy (michele_laino):

thank you! :)